OpenStudy (anonymous):

A 57-foot long piece of rope is cut into two pieces so that one piece is three feet shorter than four times the other piece. Find the length of each piec

OpenStudy (anonymous):

let the length of one piece be x so the other piece=4x-3 x+4x-3=57 5x=57+3=60 so x=60/5=12 so the other piece=45